# The generative quantum eigensolver (GQE) and its application for ground state search

### v1

@@ 3.2 The Effect of Pre-Training
We also conduct a pre-training experiment, focusing specifically on the config-to-config transfer scenario in $\texttt{N}_{2}$. Let $\hat{H}(d)$ represent the $\texttt{N}_{2}$ Hamiltonian at bond length $d$. $\hat{H}(d)$ can be expressed as a weighted sum of Pauli operators:
$\hat{H}(d)=\sum_{a=1}^{N_{H}}h_{a}(d)\hat{P}_{a},$ (8)
where $h_{a}(d)$ denotes the weights. We transfer the dataset generated during the training with $\hat{H}(1.2)$ to a dataset for pre-training with $\hat{H}(1.4)$. The initial training for $\hat{H}(1.2)$ comprises 500 steps, and at each step, we sample 50 sequences. Consequently, from the training, we compile the dataset $\mathcal{D}^{(1.4)}_{\textrm{original}}:=\{\vec{j}_{m},E^{(1.4)}(\vec{j}_{m})\}_{m=1}^{M}$, where $E^{(d)}(\vec{j})$ estimates $\textrm{Tr}\left(\hat{H}(d)U(\vec{j})\rho_{\textrm{0}}U(\vec{j})^{\dagger}\right)$ and $M=500\times 50=25,000$. From $\mathcal{D}^{(1.4)}_{\textrm{original}}$, we omit relatively high-energy data having $E^{(1.4)}(\vec{j}_{m})>-107.45$ and construct the dataset $\mathcal{D}^{(1.4)}$, which has $\sim 14,700$ data.
In the pre-training phase, $\mathcal{D}^{(1.4)}$ is divided into 294 batches. The transformer is then pre-trained using each batch once. Subsequently to this pre-training, the transformer undergoes further training for 500 steps. The hyper-parameters are set to be the same as in the experiment in Section 3.1.
In Figure 6, we display training performance following pre-training. We conducted ten trials of the experiment. In each trial, the transformer’s parameters are randomly initialized with different seeds, but the same dataset $\mathcal{D}^{(1.4)}$ is used for pre-training. Let $E_{\min}(s)$ be the minimum energy found by step $s$ in each trial. The left figure presents the mean of $E_{\min}(s)$ and its standard deviation across the ten training runs. Let $E_{\text{avg}}(s)$ be the average of energies generated at step $s$ in each trial. The right figure illustrates the mean of $E_{\text{avg}}(s)$ and its standard deviation. In each figure, the result is depicted by a green circle. As a baseline, we also include results from training conducted without pre-training. Ten trials of this experiment are conducted, which include three trials from the experiment described in Section 3.1. In each figure, the benchmark result is represented by a gray triangle.
The left figure indicates that the pre-training helps to find a lower energy in this example. The minimum energy found at the final step is $\sim 0.01$ (Hartree) lower on average in the training after pre-training than without pre-training, which corresponds to $\sim 25\%$ improvement in terms of the deviation from the exact energy ($\simeq-107.591$ Hartree). On the other hand, we observe that the average energy is almost unchanged with the pre-training, as shown in the right figure. This phenomenon can be interpreted as encouraging the pre-training to search for a wider variety of sequences; the fluctuation of the energy value around $\text{step}=50\sim 100$ in the right figure may indicate this. This diversity in sequences aids in finding lower energy configurations but also results in generating sequences with higher energy. Consequently, the average energy remains similar to that observed in training without pre-training.
@@ 4 Conclusion and Discussion
We propose the GQE algorithm, a novel method that applies a generative model to obtain quantum circuits with desired properties. In particular, we introduce GPT-QE, which is based on the transformer architecture, and we design its training and pre-training schemes using the logit-matching technique. We also introduce and discuss scenarios for obtaining datasets usable in pre-training, as well as proposals for extending the architecture’s capabilities by including inputs. In our numerical experiment, we address the problem of searching for the ground state of the electronic structure Hamiltonians of several molecules: $\texttt{H}_{2}$, LiH, $\texttt{BeH}_{2}$, and $\texttt{N}_{2}$ We demonstrate that GPT-QE finds a quantum state with energy close to that of the ground state. The efficacy of the training scheme that utilizes logit-matching is confirmed through the average deviation from random circuit generation. Additionally, we validate the effectiveness of pre-training; our results show that, by utilizing pre-training, we can train the transformer successfully without operating a quantum device and significantly reduce the total number of quantum circuit runs.
@@
Many research directions should be explored to fully enable the GPT-QE models to progress beyond proof of concept. First, it is essential to validate the performance of GPT-QE on an actual quantum device and analyze its robustness to noise. Additionally, experiments with larger molecules are required to assess the optimization behavior. Another consideration is how to effectively integrate GPT-QE with the VQE framework. A straightforward approach might involve using VQE as a post-processing step, but there are many opportunities for more integrated hybridization. As a particularly interesting existing method for hybridization, the ADAPT-VQE strategy is known to achieve high accuracy [30], though it necessitates numerous quantum circuit runs. Combining ADAPT-VQE with GPT-QE is another potential direction for future research.
@@
Each component of GPT-QE is also open to updates and improvements. In our numerical experiment, we employ a chemically inspired operator pool, but, theoretically, any kind of operators could be included. The transformer’s design could be modified to generate a sequence of tokens and the parameters embedded in quantum gates. Such an extension would facilitate easier hybridization with VQE. To fully leverage the pre-training feature, exploring how to design suitable inputs for the transformer is crucial.
@@
As mentioned earlier in the paper, applying and extending the GQE framework to problems beyond ground-state approximation is also feasible. For instance, if GQE can accept classical data inputs, it could be applied to supervised machine-learning problems. The critical question is determining which types of machine learning problems are best suited for the GQE framework. This inquiry requires careful study and identification of suitable problems.
The significance of pre-training underscores the need for data storage and sharing as common resources. We can take inspiration from the machine learning community, which shares multiple datasets for training and testing purposes. Through such efforts, we anticipate that practical quantum applications using the generative models will soon be achievable.
@@
We hope that, in a parallel track, and a decade later from when some of us introduced the VQE, the community will embrace GQE and work together to enable many of the extensions briefly proposed above to help in achieving the goal of near-term quantum utility.
@@ Acknowledgements
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0027330. K.N. acknowledges the support of Grant-in-Aid for JSPS Research Fellow 22J01501. L.B.K. acknowledges support from the Carlsberg Foundation. A.A.-G. acknowledges support from the Canada 150 Research Chairs program and CIFAR as well a the generous support of Anders G. Frøseth. We are deeply grateful to the Defense Advanced Research Projects Agency (DARPA) for their generous support and funding of this project, under the grant number HR0011-23-3-0020.
Appendix A The benchmark experiment
In Section 3.1, we show that GPT-QE successfully finds a low-energy state. However, to verify that the training correctly works, we need to compare the performance with the case where training is not involved. For this purpose, we perform a benchmark experiment that randomly samples 10 (for $\texttt{H}_{2}$) or 40 (for other molecules) tokens from $\mathcal{G}$ and constructs quantum circuits. In each trial, we generate the same number of sequences as those generated in GPT-QE. We conduct three trials.
In Figure 7, we provide the mean of the minimum energy found in each trial and its standard deviation for each bond length in each molecule. The benchmark result is shown by the gray cross mark and the other result is shown with the same symbol as in Fig. 4. This confirms that the training scheme successfully guides the transformer to find lower energy, especially for the $\texttt{N}_{2}$ molecule. This verifies the validity of the logit-matching technique.
Figure 7: The average training results with the electronic structure Hamiltonians: $\texttt{H}_{2}$ (top left), LiH (top right), $\texttt{BeH}_{2}$ (bottom left), and $\texttt{N}_{2}$ (bottom right) in sto-3g basis. The results of GPT-QE are depicted as green points (gpt-qe). The Hartree-Fock energy is indicated by a gray dotted line (hf), and the exact full configuration interaction (CI) energy, calculated through diagonalization, is represented by a black line (exact). For comparison, we also include results from randomly sampling gates from $\mathcal{G}$, which are indicated by gray cross marks (benchmark). For each data point, we run three trials and show the mean value and its deviation. When the error bar is not shown, the standard deviation is within the size of the data point.

### v2

@@ 3.2 The Effect of Pre-Training
To investigate the effectiveness of pre-training in GPT-QE, we employed the coefficient reweighting technique described in Appendix C. This approach allows us to transform 75,000 (sequence, energy) pairs obtained from 1,500 epochs of training for $\texttt{N}_{2}$ at bond length 1.2 Å into data suitable for bond length 1.05 Å. To retain only the most useful data, we selectively keep the top xx% of the lowest energy data.
Table 1 summarizes the statistics of the pre-constructed datasets used for different pre-training ratios. The datasets contain sequences with energies ranging from the minimum energy of -107.4631 Hartree to higher energy values, with the mean energy increasing as the percentage of retained data increases. Note that the FCI energy for this system is -107.5854 Hartree, indicating that the pre-constructed data consists of significantly higher energy configurations compared to the exact ground state.
Table 1: Statistics of pre-constructed datasets for different pre-training ratios.
Pre-training Ratio # of Sequences Energy (Hartree)
Min Mean Std
Top 5% 3,760 -107.4631 -107.4193 0.0060
Top 10% 7,520 -107.4631 -107.4165 0.0051
For incorporating the pre-constructed data into the training process, there are two possible approaches: (i) initially loading pre-constructed data and then starting to load model-generated data, or (ii) mixing a certain proportion of pre-constructed data into the model-generated data during training. We adopted approach (ii): mixing a certain proportion of pre-constructed data into the model-generated data during training. Specifically, we initially set 30% of the replay buffer data to be pre-constructed data, with the remaining 70% generated by the model. The proportion of pre-constructed data is then linearly decreased over 150 epochs until it reaches 0%. This approach serves as a demonstration of pre-training capabilities, and finding the optimal strategy requires further investigation.
Figure 7 shows the number of energy evaluations required to reach chemical accuracy for the $\texttt{N}_{2}$ molecule at bond length 1.05 Å as a function of the percentage $x$ of top low-energy data used for pre-training. The 0% case represents training without any pre-constructed data.
The results demonstrate that pre-constructed data significantly reduces the number of energy evaluations needed to achieve chemical accuracy. Specifically, using 5% of the lowest energy data reduces the energy evaluation count by approximately 8%, while using 10% of the data reduces it by approximately 17% compared to the baseline without pre-training. Overall, the pre-constructed data enables a reduction of nearly 20% in the energy evaluations required to reach chemical accuracy, highlighting the effectiveness of pre-training in improving the efficiency of the GPT-QE training process.
Figure 7: Effect of pre-training data ratio on energy evaluation count. The plot shows the number of energy evaluations required to reach chemical accuracy as a function of the percentage of low-energy data used for pre-training. Error bars represent standard error across multiple runs.
3.3 Demonstrations on quantum device
To validate the practical applicability of GPT-QE, we conducted experiments on the IBM 127-qubit backend ibm_kawasaki for the $\texttt{H}_{2}$ molecule. We selected four qubits (Qubits [1,2,3,4]) with a linear layout characterized by low readout error rates and low two-qubit error rates and to minimize the impact of device noise on our calculations. The expectation value of each Pauli term in the Hamiltonian was measured using 8,192 shots per measurement.
Given the inherent noise in current quantum devices, we applied readout error mitigation and Zero-Noise Extrapolation (ZNE) techniques [32], as implemented in Qiskit 1.1.1 [33], to reduce systematic errors caused by device noise. To further minimize the impact of device noise on calculation accuracy, we employed an adaptive training strategy: we initially set the number of tokens to 2 and gradually increased it until the energy values converged. For each fixed sequence length, GPT-QE was trained for 25 steps with $N_{\text{batch}}=N_{\text{sample}}=N_{\text{buffer}}=10$ and $N_{\text{iter}}=1$. The training utilized Logit Matching loss as described in Appendix B.
Figure 8 presents the results for $\texttt{H}_{2}$ molecular energies across different bond lengths. The calculations performed on ibm_kawasakiibm\_kawasaki closely reproduce the exact energy values, demonstrating successful training on the quantum device. We compared the quantum device results with statevector simulations using the same optimal sequences obtained from GPT-QE calculations. The GPT-QE energies from the quantum device lie within 0.002 Hartree of the ground-state energies from statevector simulations, confirming the effectiveness of our error mitigation approach.
We observed that the required sequence length increases with the $\texttt{H}_{2}$ bond length. Specifically, for bond lengths shorter than 0.7Å, 2 tokens suffice, while for bond lengths of 1.2Å and 2.0Å, the required number of tokens increases to 3 and 6, respectively. This trend reflects the growing importance of electron correlation effects, which Hartree-Fock theory cannot capture. GPT-QE effectively captures these correlation effects by adaptively increasing the number of tokens, demonstrating its ability to handle varying levels of electronic complexity.
Figure 8: Comparison of $\texttt{H}_{2}$ molecular energies at different bond lengths: exact diagonalization results (black line), GPT-QE calculations on the IBM quantum device with error mitigation (green spheres), and statevector simulations using optimal sequences from quantum device training (red triangles). The quantum device results closely match the exact energies, demonstrating successful training and error mitigation.
@@ 4 Conclusion and Discussion
We propose the GQE algorithm, a novel method that applies a generative model to obtain quantum circuits with desired properties. In particular, we introduce GPT-QE, which is based on the transformer architecture, and we design its training scheme. We also introduce and discuss proposals for extending the architecture’s capabilities by including inputs. In our numerical experiments, we address the problem of searching for the ground state of the electronic structure Hamiltonians of several molecules: $\texttt{H}_{2}$, LiH, $\texttt{BeH}_{2}$, and $\texttt{N}_{2}$ . We demonstrate that GPT-QE finds a quantum state with energy close to that of the ground state.
@@
Many other research directions should be explored to fully enable the GPT-QE models to progress beyond this proof of concept. First of all, conditional generalization described in Section 2.2.3 is an important direction to fully utilize the power of machine learning. Additionally, experiments with larger molecules are required to further assess the optimization behavior. Another consideration is how to effectively integrate GPT-QE with the VQE framework. A straightforward approach might involve using VQE as a post-processing step, but there are many opportunities for more integrated hybridization. As a particularly interesting existing method for hybridization, the ADAPT-VQE strategy is known to achieve high accuracy [34], though it necessitates numerous quantum circuit runs. Combining ADAPT-VQE with GPT-QE is another potential direction for future research.
@@
Each component of GPT-QE is also open to updates and improvements. In our numerical experiment, we employ a chemically inspired operator pool, but, theoretically, any kind of operators could be included. The transformer’s design could be modified to generate a sequence of tokens and the parameters embedded in quantum gates. Such an extension would facilitate easier hybridization with VQE. To fully leverage the transformer’s capabilities, exploring how to design suitable inputs for the transformer is also crucial.
@@
As mentioned earlier in the paper, applying and extending the GQE framework to problems beyond ground-state approximation is also feasible. For instance, Ref. [18] has already investigated the application of a GQE model with classical inputs to the problem of binary optimization. One could also envision applying such input-conditioned GQE models to supervised machine-learning problems. The critical question would then be determining which types of machine learning problems are best suited for the GQE framework. This inquiry requires careful study and identification of suitable problems.
@@
We hope that, in a parallel track, and a decade later from when some of us introduced VQE, the community will embrace GQE and work together to enable many of the extensions briefly proposed above to help in achieving the goal of near-term quantum utility.
@@ Acknowledgements
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0027330. K.N. acknowledges the support of Grant-in-Aid for JSPS Research Fellow 22J01501. L.B.K. acknowledges support from the Carlsberg Foundation. A.A.-G. acknowledges support from the Canada 150 Research Chairs program and CIFAR as well as the generous support of Anders G. Frøseth. We are deeply grateful to the Defense Advanced Research Projects Agency (DARPA) for their generous support and funding of this project, under the grant number HR0011-23-3-0020. A part of this work was performed for Council for Science, Technology and Innovation (CSTI), Cross-ministerial Strategic Innovation Promotion Program (SIP), “Promoting the application of advanced quantum technology platforms to social issues” (Funding agency: QST).
We acknowledge that nearly simultaneously with the second version update of this work, [35] proposed improvements to our original GQE learning approach by incorporating mechanisms equivalent to modified loss functions and replay buffers.
Appendix A Details of the experiment
A.1 Molecule Generation
The molecular systems are generated using the CUDA Quantum solver [31] with the following configuration:
1import cudaq_solvers as solvers
2
3molecule = solvers.create_molecule(geometry, molecule.basis, spin=0, charge=0, nele_cas=nele_cas, norb_cas=norb_cas, ccsd=True, casci=True)
where geometry is the geometry of the molecule, molecule.basis is the basis set, nele_cas is the number of electrons in the active space, and norb_cas is the number of orbitals in the active space. The molecule object generation includes the calculation of the Hartree-Fock energy, the CCSD energy, and the FCI energy based on PySCF [36].
A.2 Operator Pool Configuration
Our operator pool consists of Pauli time evolutions plus the identity operator: $\mathcal{G}=\{e^{i\hat{P}_{j}t_{j}}\}_{j}\cup\{I\}$, where $\hat{P}_{j}$ represents a tensor product of Pauli operators, $t_{j}$ is a real number, and $I$ is the identity operator. Defining $\mathcal{P}$ as the set of $\hat{P}_{j}$ and $\mathcal{T}$ as the set of $t_{j}$, the size of the operator pool is $|\mathcal{P}|\times|\mathcal{T}|+1$.
For $\mathcal{P}$, we derive chemically inspired choices from the unitary coupled-cluster single and double excitations (UCCSD). Letting $T$ denote the sum of all fermionic excitation operators included in UCCSD, $\mathcal{P}$ is selected such that $e^{iP_{\ell}\theta_{\ell}}(P_{\ell}\in\mathcal{P})$, with an angle $\theta_{\ell}$, is part of the decomposed operators when $e^{T-T^{\dagger}}$ is broken down into Pauli time evolutions by the Trotter decomposition [37]. For $\mathcal{T}$, we choose $\mathcal{T}=\left\{\pm 2^{k}/320\right\}_{k=0}^{5}$. In our implementation, we use the CUDA Quantum solver [31] to generate the operator pool:
1import cudaq_solvers as solvers
2import numpy as np
3
4pool = []
5params = [2**k/320 for k in range(0, 5)] + [-2**k/320 for k in range(0, 5)]
6operators = solvers.get_operator_pool("uccsd", n_qubits, n_electrons)
7pool.append(get_identity(n_qubits))
8for o in operators:
9 for p in params:
10 pool.append(exp_pauli(p * o)) # Create a scaled Pauli operator
where n_qubits is the number of qubits in the quantum system and n_electrons is the number of electrons in the molecule. Here, get_identity(n_qubits) returns the identity operator for the quantum system, and exp_pauli(P) implements the exponential operator $e^{iA}$ for a given operator $A$.
A.3 Adaptive scheduling of inverse temperature $\beta$
In the GPT-QE setup, the inverse temperature $\beta$ governs the exploration-exploitation trade-off. We employ a dispersion-triggered schedule that adapts $\beta$ based on the variance of the energies generated at each training iteration. Let $\{E(\vec{j}_{m})\}_{m=1}^{M}$ be the energies evaluated at iteration $t$ and define
$\beta_{t+1}=\begin{cases}\beta_{t}-\alpha,&\text{if }\operatorname{std}\!\left(\{E(\vec{j}_{m})\}_{m=1}^{M}\right)<\tau_{\mathrm{disp}},\\[2.0pt] \beta_{t}+\alpha,&\text{otherwise},\end{cases}$ (4)
with a fixed step size $\alpha=0.02$ and dispersion threshold $\tau_{\mathrm{disp}}=10^{-5}$ in all experiments. Intuitively, when the batch collapses (very small dispersion), we decrease $\beta$ to re-broaden the sampling distribution and preserve exploration; otherwise, we increase $\beta$ to gradually sharpen the distribution and intensify exploitation around lower-energy regions discovered so far. Compared to a monotonic temperature-increase schedule, this dispersion-triggered update is less prone to entrapment in local minima and empirically improves convergence. Note that this strategy utilizes the fact that the model only needs to sample the optimal circuit a high fraction of the time, but does not need to converge to a collapsed distribution sampling only the optimal circuit.
Appendix B Loss Functions
B.1 Logit Matching Loss
The logit-matching loss is designed to align the model’s cumulative logits $w_{\mathrm{sum}}(\vec{j}_{m};\theta)$ with the evaluated energies $E(\vec{j}_{m})$, inducing a pseudo-thermal preference for lower-energy circuits. This approach encourages the model to generate circuits with lower energies by making the logits correspond to the energy landscape.
Logit matching is inspired by the following argument. Let us consider the process sampling $U_{N}(\vec{j})$ according to $p_{N}(\beta,\vec{j})$ and applying it to $\rho_{0}$. Let $\rho(\beta)$ be the quantum state generated by the stochastic process. With $\mathcal{E}_{N}(\vec{j},\rho):=U_{N}(\vec{j})\rho U_{N}(\vec{j})^{\dagger}$, it can be written as
We observe that if $w_{\textrm{sum}}(\vec{j})=E(\vec{j})$ is satisfied, $\rho(\beta)$ gives a pseudo thermal state with the inverse temperature $\beta$ in the sense that the quantum state is generated according to the probability $\exp\left(-\beta E(\vec{j})\right)$. Therefore, increasing the value of $\beta$ creates a bias towards generating lower energy quantum states.
From these observations, we design the logit matching loss so that $w_{\textrm{sum}}(\vec{j})\simeq E(\vec{j})$ is satisfied. The loss function is defined as
$\mathcal{L}_{\textrm{LM}}=\frac{1}{N_{\textrm{batch}}}\sum_{m=1}^{N_{\textrm{batch}}}\left(e^{-\beta w_{\textrm{sum}}(\vec{j}_{m})}-e^{-\beta E(\vec{j}_{m})}\right)^{2}.$ (6)
For numerical stability, we add an offset to the output of the quantum device. More specifically, when calculating the cost function above, we substitute $E(\vec{j})+E_{\textrm{offset}}$ instead of the original $E(\vec{j})$. The value of $E_{\textrm{offset}}$ is chosen to be 107 for $\texttt{N}_{2}$, in the experiment for Fig. 6. Since we match the sum of logits with the energy function, we call this technique logit-matching.
B.2 GRPO-based Loss Function
While the logit-matching loss (Eq. (6)) encourages a pseudo-thermal preference, it does not explicitly minimize the expected energy; empirically, this can lead to periods where the loss decreases yet the energy plateaus above the ground-state level. To remedy this, an energy-oriented policy surrogate is added to amplify the probability of lower-energy samples within each iteration. Specifically, we instantiate this surrogate using the Group Relative Policy Optimization (GRPO) [26] objective, which computes so-called advantages by normalizing rewards within a group of sampled sequences. For a sequence $\vec{j}_{m}=\{j_{m,k}\}_{k=1}^{N}$, define the sequence-level reward and advantage
$r_{m}:=-E(\vec{j}_{m}),\qquad\hat{A}_{m}:=\frac{r_{m}-\operatorname{mean}(\{r_{m}\}_{m=1}^{M})}{\operatorname{std}(\{r_{m}\}_{m=1}^{M})},$ (7)
which yields variance reduction and scale invariance across iterations. We further define the importance ratio
$\rho_{m,k}\;:=\;\frac{\pi_{\theta}(j_{m,k}\mid q,j_{m,<k})}{\pi_{\theta_{\mathrm{old}}}(j_{m,k}\mid q,j_{m,<k})}.$ (8)
The function $\pi_{\theta}(j_{m,k}\mid q,j_{m,<k})$ is the probability function that $j_{m,k}$ is generated as the $k$-th token given the k1k-1 previously generated tokens $j_{m,<k}$ and a fixed start token, and it corresponds to $\exp\left(-\beta w^{(k)}_{j_{m,k}}\right)$ in the notation introduced in Section 2.2.1. The function has the parameters $\theta$ included in the neural network written out explicitly. The symbol $\theta_{\mathrm{old}}$ denotes frozen reference parameters, taken as the initial parameters in all of our experiments. Using $\rho_{m,k}$, the individual GRPO-based loss function for each batch is defined as
$\mathcal{L}_{\text{GRPO}}=\frac{1}{N_{\textrm{batch}}}\sum_{m=1}^{N_{\textrm{batch}}}\frac{1}{N}\sum_{k=1}^{N}\mathrm{clip}\left(\rho_{m,k},\,1-\varepsilon,\,1+\varepsilon\right)\,\hat{A}_{m},$ (9)
[…]
Appendix C Coefficient reweight to Increase the Dataset
In the context of quantum computation, the electronic structure Hamiltonian can be mapped to a linear combination of Pauli operators:
$\hat{H}(\vec{\Delta})=\sum_{a=1}^{N_{H}}h_{a}(\vec{\Delta})\hat{P}_{a},$ (10)
where $h_{a}(\vec{\Delta})$ are real coefficients when the geometry is $\vec{\Delta}$ and $\hat{P}_{a}$ are tensor products of Pauli operators. When we evaluate the expectation value, we need to measure each individual Pauli expectation value $\langle\hat{P}_{a}\rangle$. Therefore, in the training of GPT-QE with $\hat{H}(\vec{\Delta})$, we obtain the dataset $\{\vec{j}_{m},q_{a}(\vec{j}_{m})\}_{m,a}$ as well as $\{\vec{j}_{m},E(\vec{\Delta},\vec{j}_{m})\}_{m}$, where $q_{a}(\vec{j}_{m})$ is the estimated value of $\langle\hat{P}_{a}\rangle$ when the circuit is constructed with the sequence of tokens $\vec{j}_{m}$ and $E(\vec{\Delta},\vec{j}_{m})=\sum_{a}h_{a}(\vec{\Delta})q_{a}(\vec{j}_{m})$.
The estimated values ${q_{a}(\vec{j_{m}})}_{m,a}$ can be used to construct $\{\vec{j}_{m},E(\vec{\Delta}^{\prime},\vec{j}_{m})\}_{m}$ with a different geometry $\Delta^{\prime}$ by simply combining the measured Pauli expectation values using different coefficients $h_{a}(\vec{\Delta}^{\prime})$. It can then be used for pre-training when solving the ground-state search of the Hamiltonian $H(\vec{\Delta}^{\prime})$.